# Cross-Entropy Games for Language Models: From Implicit Knowledge to General Capability Measures

### v1

@@ 4.2.3. Fine-Tuning Definition
The other, somehow deeper and more general, notion of playability pertains to the ability of models to get better by fine-tuning performed for long enough. It should be noted that closed models are not always available in fine-tuning mode, and so that this notion is strictly less applicable than the few-shot playability one:
@@
A well-posed game $G$ is playable in fine-tuning mode by a model $\mathcal{M}$ if for some time $\mathcal{T}>0$, the model $\mathcal{F}_{G}^{\mathcal{T}}\left(\mathcal{M}\right)$ is strictly better in expectation than the model $\mathcal{M}$ at $G$.
@@
It is generally expected that if a game is playable in fine-tuning mode, it is also playable in zero-shot mode.
As we will see in Section 4.3, the fine-tuning playability is crucial for us to relate games with one another via the so-called transfer value. This notion of transfer value then allows us to build a geometric structure on the space of Xent Games, which we rely upon to construct relevant Xent Measures.
@@ 4.3. Transfer Value
As discussed in Section 4.2, a playable game is one where a player’s strategy can be adjusted by experience, leading to a higher score; in other words, the model $\mathcal{M}$ is learning something. But is it worth playing the game, i.e. is the learned information valuable? Surely, this is in large part very subjective. In this subsection, we will define the value in terms of the increased abilities of an agent fine-tuned to play that game. This of course shifts the subjectivity further: what is the end-goal in terms of game playing utility? We defer the investigation of the latter problem to Section 5.
@@ 4.3.1. Transfer Value of a Game
The transfer value should simply be understood as ‘how much value does it bring towards playing $G_{2}$ to play $G_{1}$?’ (normalized by for how long one plays $G_{1}$). Note that there is no reason to expect symmetry, i.e. $\mathcal{V}_{G_{1}}\left(G_{2}\right)$ is generally not equal to $\mathcal{V}_{G_{2}}\left(G_{1}\right)$.
@@
As much as we will often take for granted that $\mathcal{M},\mathcal{F}$, and $\mathcal{T}$ have been suitably set, working with transfer values, it is important to remind ourselves that $\mathcal{V}$ depends a priori on them, and that a ‘good setting’ for $\mathcal{M},\mathcal{F},\mathcal{T}$ has been found. It is however important to keep in mind that we can still expect that the final outputs of our constructions (see Sections 5 and 6 below) to not depend on $\mathcal{M},\mathcal{F},\mathcal{T}$: we expect all ‘good’ $\mathcal{M},\mathcal{F},\mathcal{T}$ settings give roughly the same general ability measures (see Section 6.2.3 for a discussion of this universality idea).
@@
While the transfer value of a game of course depends on the many details abstracted in the $\mathcal{F}$ mechanism (and of the game metadata, in particular the models linked to the judge and players), it is an interesting question to ask how much the constructions we build from it depend on such details (see Section 6).
@@ 4.3.2. Transfer Value: Numerical Example
The transfer value concepts introduced above can be illustrated on a simple triplet of games (see also Section 3.1 for similar games).
@@
Experimentally, we show the presence of transfer learning (i.e. nontrivial transfer value) by making three independent fine-tuning (one on each of the three games) of a model (Qwen3 8B in FP8 in our experiment with GPT2-XL as the judge model): for each of the three models being fine-tuned, we evaluate their performance on all three games. The fine-tuning procedure $\mathcal{F}$ is performed using the REINFORCE algorithm [Wil92] on rewards associated with the choices of $\mathcal{M}$ at every $elicit$ call, run with some the AdamW optimizer and learning rate $10^{-6}$ with a token processing budget $\mathcal{T}$, with a corpus of 20k ‘stories’; the evaluation is performed on a corpus of 600 ‘stories’. Our experiments demonstrate a clearly non-trivial transfer value across games.
@@ 5. Xent Game Measures: Scopes and Covers
In this section, we construct, Xent Measures, which yield theoretically-motivated LLM benchmarks from Xent Game scopes, which are families of playable Xent Games representing a desirable set of abilities. This will pave the way towards Section 6, which defines general capability scopes, leading to general ability measures.
@@ 5.1.1. From specific to general abilities: benchmarking
Traditionally, the benchmarking of LLMs tests their explicit knowledge using a set of questions (that are typically made by humans) with e.g. multiple-choice answers. Designing such benchmarks is of course very contingent on specific areas; there is also naturally the risk of data contamination. See e.g. [LBL23, PCDSC24, CWWWX23] for surveys on the works.
Interesting examples of benchmarks that attempt to go towards more general abilities include for instance [Cho19, CKKLP25]. A very promising line of work that generally attempts at evaluating LLM performance is via the creation of environments to evaluate them, see e.g. [XDCGZ24, SSSJ25, CKATT18], though a principled way to extract a benchmark from such environment is not yet available. The idea to use games has been popularized in e.g. [CZJGS24].
In spite of there being a plethora of mechanisms to benchmark LLMs, the interpretations and theoretical underpinnings of such benchmarks are arguably rarely clear. In fact, one could even say that it is not very clear what one expects from a benchmark, besides answering the informal (and very underspecified) question of which LLMs are ‘better, more capable’. Since these notions are not very well defined a priori, if one asks the question ‘what is expected from a new benchmark’, a key element one expects to inform us along some dimension that is not captured accurately by other benchmarks, but that one will recognize (perhaps a posteriori) as being driven by useful skills.
While the practical successes yielded in part by the current state of the art is undeniable (given that, ultimately, most are able to recognize spectacular abilities of LLMs, without usually being able to define them), this state of affairs makes it naturally hard to ascribe a specific meaning to a benchmark score or to clearly identify needs for future benchmark scores.
@@ 5.2. Score Normalization
As discussed above, the goal of benchmarking is naturally to compare various models (whether of different architectures, datasets or simply fine-tuning of a given model). Naturally, each playable Xent Game $G$ instance can be used to measure various models; given the diversity of such games, the question of score normalization, in particular across games, roles, and seeds, is not obvious.
@@ 5.2.1. Base Model and Few-Shot Mode
Related to score normalization is that of units: what are the units in which we aim to understand the performance of models playing Xent Games? While the Xent Games have the natural advantage of having scores expressed in bits, surely the bits are not each of the same value; even within the family of games that are playable for some model, some are much harder to play than others.
For instance, going from 0.95 win rate at playing chess against an agent to 0.999 win rate could be interpreted as a very drastic increase in ability, while it would correspond to a few percent of score increase (much more than going from a rate of 0.5 to 0.55, say).
In order to normalize scores, a simple solution is to identify a sufficiently strong base model $\mathcal{M}_{\mathcal{B}}$, find reasonable values of $N$ (number of samples), $\epsilon>0$ (tolerated ‘failure probability’), $\theta$ (increment value) and $m\geqslant 2$ (number of increments), and to consider the set of $\left(N,\epsilon,\theta,m\right)$-playable games relative to $\mathcal{M}_{\mathcal{B}}$ (see Definition 4.2.1), which we will simply call (unless specified otherwise) the playable Xent Games.
@@ 5.2.2. Normalized Score Definition
  • If $\tau=\ell$ (i.e. $\mathcal{M}$ has not outpaced $\mathcal{M}_{\mathcal{B}}$ in $\ell$ steps) and $arms_{N,\tau}=b+j\alpha$ for some $j\leq m$, then $narms_{N}$ is exactly $j$; anything in between is the unique affine interpolation that is continuous in with respect to $arms_{N,\tau}$.
@@ 5.2.3. Multi-Game Benchmark Design
In Section 5.2, we presented the question that a unified normalization aims to answer the following: What is the performance of a model playing a game, relative to a base model?
@@
The problem of selection of the games appears however to be substantially more subjective and challenging: we discuss this question in Section 5, and an answer based on an evolution-inspired meta-game will be provided in Section 6.
@@ 5.3. Benchmarks from Xent Measures
As a way to propose a solution to the benchmarking problem raised in Section 5.1 above, from to the structure of the Xent Game space 2.3.3 it is tempting to suggest to rely on benchmarks of the following forms for a suitable (finite) family of games $\Gamma$:
@@
  • The covering is ‘fair’, i.e. there is no ‘over-representation’ of some games on the scope.
@@ 5.4. Xent Measure from Scope
The idea behind the definition of Xent Measures from scopes is simple, and very related to meta-learning (see e.g. [Bax00] for foundations, [SYSKZ24] for a recent survey about LLMs, and [BVLFW25] for a survey about meta-reinforcement learning): given a scope $\Sigma$, find a minimal covering family $\Gamma$ such that any game in $G\in\Sigma$ is ‘close enough’ to a game in $G^{\prime}\in\Gamma$ (in the sense that $\mathcal{V}_{G^{\prime}}\left(G\right)$ is large enough): in other words, we want a minimal set of games $\Gamma$ such that if we had to play any game $G\in\Sigma$, and if we are able to play the games in $\Gamma$, we could quickly fine-tune our model to play $G$.
A minimal covering family $\Gamma$ naturally yields a capability benchmark for LLMs on the scope $\Sigma$: for a model $\mathcal{M}$, consider as the (histogram) of the scores $S_{G}\left(\mathcal{M}\right)$ for $G\in\Gamma$. Intuitively, the idea is that if $\Sigma$ is dense, each $G\in\Gamma$ plays a similar role, and can thus be considered interchangeably, with the minimality guaranteeing ‘no over-representation’ of certain regions of the game space: the histogram of the scores then gives us a good idea of how well $\mathcal{M}$ can be expected to perform on a random game sampled ‘near $\Sigma$’ (i.e. in a region well covered by $\Sigma$).
@@
  • For a family of Xent Games $\Gamma\subset\Upsilon$, we denote by $\Delta_{\Gamma}\left(\Sigma\right)$ the transfer cover of $\Sigma$, defined as $\min_{G\in\Sigma}\max_{G^{\prime}\in\Gamma}\mathcal{V}_{G^{\prime}}\left(G\right)$ : this represents how much $\Sigma$ is guaranteed to be covered by $\Gamma$.
@@
We have that $\Gamma_{\Sigma}^{\eta}$ is in some sense a ‘skeleton’ of $\Sigma$, i.e. a version of $\Sigma$ stripped of duplicate games, and possibly games replaced by ‘leaner versions of themselves, for games $G$ such that there exists a $G^{\prime}$ with $\mathcal{V}_{G^{\prime}}\left(G\right)>\mathcal{V}_{G}\left(G\right)$.
@@
For a scope $\Sigma$, and a cover $\eta$ we call benchmark measure associated with $\left(\Sigma,\eta\right)$ the Xent Measure $\mu_{\Sigma}^{\eta}:=\mu_{\Gamma_{\Sigma}^{\eta}}$.
@@
Even if the scope $\Sigma$ is large, most of the interesting covers should correspond to relatively small values $\eta$: these correspond to fairly coarse covers.
@@ 5.5. Scope Selection
These considerations can apply separately or simultaneously: for instance, we may have a set of pre-defined goals, but would like to expand in generality around those, or may want to set goals to allow a model to differentiate itself from other models.
Since the key problem raised by the Items 1 and 2 is to construct a specific list of games (whether hand-crafted or picked according to a feature selection process), we see that in any case the problem is the following: to answer general abilities question 3, starting from a list of specific games (see Problem 49) below.
@@ 5.5.1. Towards a General Ability Measure
Giving up on the idea of finite set of games appears to be a promising direction, bringing us closer to something like a curriculum. Naturally, this forgoes the idea of a uniform benchmark measure (see Remark 50 below), and leads us to focus on an order for games that are to make up a benchmark.
@@
While this question is naturally very under-specified, we will see in Section 6 that evolution-inspired ideas can lead us towards answers to the Problem 49.
@@ 6. Evolution in Game Space
In Section 5, we introduced the idea of benchmarks based on scopes, i.e. families of Xent Games. In Section 5.5.1, we discussed the challenges associated with infinite scores, and formulated Problem 49: how to find, starting from an initial family of games $\Sigma_{0}$ a sequence of measures that can be viewed (thinking of the $k\to\infty$ limit) as an approximation of general abilities (in the sense of Remark 51)?
@@ 6.1. Xent Game Space Geometry
One of the key insights about the Xent Game space is that the games are related via their structures to one another: this is essentially a consequence of the proof of Theorem 25 above. This suggests (informally) that if one has a strong enough base model $\mathcal{M}_{\mathcal{B}}$, one should be able to transform any (reasonable) Xent Game into any other Xent Game by a sequence of moves such that we can transfer some non-trivial amount of skill from game to game (locally, not necessarily across the whole path).
@@
For $\alpha>0$ and a collection of games $\Sigma_{0}$, we denote by $\left(\Sigma_{0}\right)_{\alpha}$ the $\alpha$-downstream component defined as the set of games $G$ reached by an $\alpha$-downstream path that starts from a game in $\Sigma_{0}$; similarly, we denote by $\left(\Sigma_{0}\right)^{\alpha}$ the $\alpha$-upstream component defined as the set of games reached by an $\alpha$-upstream starting from a game in $\Sigma_{0}$.
@@
For a ‘sophisticated enough model’, it is reasonable (by the construction of Theorem 25 above) to postulate that all the (playable) Xent Games are related, in the following sense: one can always modify a playable game into another one by a chain that is $\alpha$-downstream and by a chain that is $\alpha$-upstream, for some small enough $\alpha$. This is the content of the following:
@@
This may suggest to take $\alpha$ small enough; or we could also instead decide to focus on an $\alpha$-connected component for some a priori fixed $\Sigma_{0}$ (if we don’t care about general abilities, but simply general abilities around some scope).

### v2

@@ 4.2.3. Fine-Tuning Definition
The other, somehow deeper and more general, notion of playability pertains to the ability of models to get better via fine-tuning. It should be noted that closed models are not always available in fine-tuning mode, and so this notion is strictly less applicable than the few-shot playability one:
@@
A well-posed game $G$ is playable in fine-tuning mode by a model $\mathcal{M}$ if, for some time $\mathcal{T}>0$, the model $\mathcal{F}_{G}^{\mathcal{T}}\left(\mathcal{M}\right)$ is strictly better in expectation at $G$ than the model $\mathcal{M}$.
@@
It is generally expected that if a game is playable in fine-tuning mode, it is also playable in few-shot mode.
As we will see in Section 4.3, fine-tuning playability is crucial for us to be able to relate games with one another via the so-called transfer value. This notion of transfer value allows us to build a geometric structure on the space of Xent Games, which we rely upon to construct relevant Xent Measures.
@@ 4.3. Transfer Value
As discussed in Section 4.2, a playable game is one where a player’s strategy can be adjusted by experience, leading to a higher score; in other words, the model $\mathcal{M}$ is learning something. But is it worth playing the game, i.e. is the learned information valuable? Surely, this is, in large part, very subjective. In this subsection, we will define the value of learning to play a game in terms of the increased abilities of an agent fine-tuned to play that game. This, of course, shifts the subjectivity further: what is the end-goal in terms of game playing utility? We defer the investigation of this latter problem to Section 5.
@@ 4.3.1. Transfer Value of a Game
The transfer value should simply be understood as ‘how much value does playing $G_{1}$ bring towards playing $G_{2}$?’ (normalized by how long one plays $G_{1}$). Note that there is no reason to expect symmetry, i.e. $\mathcal{V}_{G_{1}}\left(G_{2}\right)$ is generally not equal to $\mathcal{V}_{G_{2}}\left(G_{1}\right)$.
@@
As much as we will often take for granted that $\mathcal{M},\mathcal{F}$, and $\mathcal{T}$ have been suitably set, working with transfer values, it is important to remind ourselves that $\mathcal{V}$ depends a priori on them, and that a ‘good setting’ for $\mathcal{M},\mathcal{F},\mathcal{T}$ has been found. It is however important to keep in mind that we can still expect that the final outputs of our constructions (see Sections 5 and 6 below) to not depend on $\mathcal{M},\mathcal{F},\mathcal{T}$: we expect that all ‘good’ $\mathcal{M},\mathcal{F},\mathcal{T}$ settings give roughly the same general ability measures (see Section 6.2.3 for a discussion of this universality idea).
@@
While the transfer value of a game depends on the many details abstracted in the $\mathcal{F}$ mechanism (and of the game metadata, in particular the models linked to the judge and players), it is an interesting question to ask how much the constructions we build from the transfer value of a game depend on such details (see Section 6).
@@ 4.3.2. Transfer Value: Numerical Example
The transfer value concepts introduced above can be illustrated on a simple triplet of games, shown in Figure 4.2 (see also Section 3.1 for similar games).
@@
Experimentally, we show the presence of transfer learning (i.e. nontrivial transfer value) by performing three independent fine-tunings (one on each of the three games) of a model (Qwen3 8B in FP8 in our experiment with GPT2-XL as the judge model): for each of the three models being fine-tuned, we evaluate their performance on all three games, as displayed in Figure 4.3. The fine-tuning procedure $\mathcal{F}$ is performed using the REINFORCE algorithm [Wil92] on rewards associated with the choices of $\mathcal{M}$ at every $elicit$ call, run with the AdamW optimizer and a learning rate of $10^{-6}$ with a token processing budget $\mathcal{T}$, with a corpus of 20k ‘stories’; the evaluation is performed on a corpus of 600 ‘stories’. Our experiments demonstrate a clearly non-trivial transfer value across games. Our results are obtained from the code available in the xega repository on GitHub.
@@ 5. Xent Game Measures: Scopes and Covers
In this section, we construct Xent Measures, which yield theoretically-motivated LLM benchmarks from Xent Game scopes, which are families of playable Xent Games representing a desirable set of abilities. This will pave the way towards Section 6, which defines general capability scopes, leading to general ability measures.
@@ 5.1.1. From specific to general abilities: benchmarking
Traditionally, the benchmarking of LLMs tests their explicit knowledge using a set of questions (that are typically made by humans) with e.g. multiple-choice answers. Designing such benchmarks is of course very contingent on specific areas; there is also naturally the risk of data contamination. See e.g. [LBL23, PCDSC24, CWWWX23] for surveys on relevant work.
Interesting examples of benchmarks that attempt to go towards more general abilities include for instance [Cho19, CKKLP25]. A very promising line of work that generally attempts at evaluating LLM performance is via the creation of environments to evaluate them in, see e.g. [XDCGZ24, SSSJ25, CKATT18], though a principled way to extract a benchmark from such environment is not yet available. The idea to use games has been popularized in e.g. [CZJGS24].
In spite of there being a plethora of mechanisms to benchmark LLMs, the interpretations and theoretical underpinnings of such benchmarks are arguably rarely clear. In fact, one could even say that it is not very clear what one expects from a benchmark, besides answering the informal (and very underspecified) question of which LLMs are ‘better and more capable’. Since these notions are not very well defined a priori, it is useful to consider the question ‘what is expected from a new benchmark’. A key element one expects of a new benchmark is for it to inform us along some dimension that is not captured accurately by other benchmarks, but that one will recognize (perhaps a posteriori) as being relevant to useful skills.
While the practical successes yielded in part by the current state of the art are undeniable (given that, ultimately, most are able to recognize spectacular abilities of LLMs, without usually being able to define them), this state of affairs makes it naturally hard to ascribe a specific meaning to a benchmark score or to clearly identify needs for future benchmark scores.
@@ 5.2. Score Normalization
As discussed above, the goal of benchmarking is naturally to compare various models (whether of different architectures, datasets, or simply fine-tuning of a given model). Naturally, each playable instance of a Xent Game $G$ can be used to measure various models; given the diversity of such games, the question of score normalization, in particular across games, roles, and seeds, is not obvious.
@@ 5.2.1. Base Model and Few-Shot Mode
Related to the question of score normalization is that of units: what are the units in which we aim to understand the performance of models playing Xent Games? While Xent Games have the natural advantage of having scores expressed in bits, surely the bits are not each of the same value; even within the family of games that are playable for some model, some are much harder to play than others.
For instance, going from a 0.95 win rate playing chess against an agent to a 0.999 win rate could be interpreted as a very drastic increase in ability (much more than going from a rate of 0.5 to 0.55, say), while it would correspond a small percent of score increase.
In order to normalize scores, a simple solution is to identify a sufficiently strong base model $\mathcal{M}_{\mathcal{B}}$, find reasonable values of $N$ (number of samples), $\epsilon>0$ (tolerated ‘failure probability’), $\theta$ (increment value), and $m\geqslant 2$ (number of increments), and to consider the set of $\left(N,\epsilon,\theta,m\right)$-playable games relative to $\mathcal{M}_{\mathcal{B}}$ (see Definition 4.2.1), which we will simply call (unless specified otherwise) the playable Xent Games.
@@ 5.2.2. Normalized Score Definition
  • If $\tau=\ell$ (i.e. $\mathcal{M}$ has not outpaced $\mathcal{M}_{\mathcal{B}}$ in $\ell$ steps) and $arms_{N,\tau}=b+j\alpha$ for some $j\leq m$, then $narms_{N}$ is exactly $j$; anything in between is the unique affine interpolation that is continuous in $arms_{N,\tau}$.
@@ 5.2.3. Multi-Game Benchmark Design
In Section 5.2, we presented the question that a unified normalization aims to answer: What is the performance of a model playing a game, relative to a base model?
@@
The problem of selection of the games appears, however, to be substantially more subjective and challenging: we discuss this question in Section 5, and an answer based on an evolution-inspired meta-game will be provided in Section 6.
@@ 5.3. Benchmarks from Xent Measures
As a way to propose a solution to the benchmarking problem raised in Section 5.1 above, it is tempting to use the structure of the Xent Game space 2.3.3 and suggest reliance on benchmarks of the following forms for a suitable (finite) family of games $\Gamma$:
@@
  • The covering is ‘fair’, i.e. there is no ‘over-representation’ of some games in the scope.
@@ 5.4. Xent Measure from Scope
The idea behind the definition of Xent Measures from scopes is simple and very related to meta-learning (see e.g. [Bax00] for foundations, [SYSKZ24] for a recent survey about LLMs, and [BVLFW25] for a survey about meta-reinforcement learning): given a scope $\Sigma$, find a minimal covering family $\Gamma$ such that any game in $G\in\Sigma$ is ‘close enough’ to a game in $G^{\prime}\in\Gamma$ (in the sense that $\mathcal{V}_{G^{\prime}}\left(G\right)$ is large enough). In other words, we want a minimal set of games $\Gamma$ such that if we had to play any game $G\in\Sigma$, and if we are able to play the games in $\Gamma$, we could quickly fine-tune our model to play $G$.
A minimal covering family $\Gamma$ naturally yields a capability benchmark for LLMs on the scope $\Sigma$: for a model $\mathcal{M}$, consider the histogram of the scores $S_{G}\left(\mathcal{M}\right)$ for $G\in\Gamma$. Intuitively, the idea is that if $\Sigma$ is dense, each $G\in\Gamma$ plays a similar role, and can thus be considered interchangeably, with the minimality guaranteeing ‘no over-representation’ of certain regions of the game space: the histogram of the scores then gives us a good idea of how well $\mathcal{M}$ can be expected to perform on a random game sampled ‘near $\Sigma$’ (i.e. in a region well covered by $\Sigma$).
@@
  • For a family $\Gamma\subset\Upsilon$, we denote by $\Delta_{\Gamma}\left(\Sigma\right)$ the transfer cover of $\Sigma$, defined as $\min_{G\in\Sigma}\max_{G^{\prime}\in\Gamma}\mathcal{V}_{G^{\prime}}\left(G\right)$: this represents how much $\Sigma$ is guaranteed to be covered by $\Gamma$.
@@
We have that $\Gamma_{\Sigma}^{\eta}$ is, in some sense, a ‘skeleton’ of $\Sigma$, i.e. a version of $\Sigma$ stripped of duplicate games, and possibly games replaced by ‘leaner’ versions of themselves, for games $G$ such that there exists a $G^{\prime}$ with $\mathcal{V}_{G^{\prime}}\left(G\right)>\mathcal{V}_{G}\left(G\right)$.
@@
For a scope $\Sigma$, and a cover $\eta$ we call the benchmark measure associated with $\left(\Sigma,\eta\right)$ the Xent Measure $\mu_{\Sigma}^{\eta}:=\mu_{\Gamma_{\Sigma}^{\eta}}$.
@@
Even if the scope $\Sigma$ is large, most of the interesting covers should correspond to relatively small values of $\eta$: these correspond to fairly coarse covers.
@@ 5.5. Scope Selection
These considerations can apply separately or simultaneously: for instance, we may have a set of pre-defined goals, but would like to expand in generality around those, or may want to set specific goals to allow a model to differentiate itself from other models.
Since the key problem raised by the Items 1 and 2 is to construct a specific list of games (whether hand-crafted or picked according to a feature selection process), if we want to complete the above picture with the generality dimension of Item 3, the following question emerges: how to grow, starting from a list of specific games, a sequence of games that measure more and more general capabilities around that initial list? This is the content of Problem 49 below.
@@ 5.5.1. Towards a General Ability Measure
Giving up on the idea of a finite set of games appears to be a promising direction, bringing us closer to something like a curriculum. Naturally, this forgoes the idea of a uniform benchmark measure (see Remark 50 below), and leads us to focus on an order for games that are to make up a benchmark.
@@
While this question is naturally very under-specified, we will see in Section 6 that ideas inspired by evolution ideas can lead us towards principled answers to Problem 49.
@@ 6. Evolution in Game Space
In Section 5, we introduced the idea of benchmarks based on scopes, i.e. families of Xent Games. In Section 5.5.1, we discussed the challenges associated with infinite scores, and formulated Problem 49: how to find, starting from an initial family of games $\Sigma_{0}$, a sequence of measures that can be viewed (thinking of the $k\to\infty$ limit) as an approximation of general abilities (in the sense of Remark 51)?
@@ 6.1. Xent Game Space Geometry
One of the key insights about the Xent Game space is that the games are related to one another via their structures: this is essentially a consequence of the proof of Theorem 25 above. This suggests (informally) that if one has a strong enough base model $\mathcal{M}_{\mathcal{B}}$, one should be able to transform any (reasonable) Xent Game into any other Xent Game by a sequence of moves such that we can transfer some non-trivial amount of skill from game to game (locally, not necessarily across the whole path).
@@
For $\alpha>0$ and a collection of games $\Sigma_{0}$, we denote by $\left(\Sigma_{0}\right)_{\alpha}$ the $\alpha$-downstream component defined as the set of games $G$ reached by an $\alpha$-downstream path that starts from a game in $\Sigma_{0}$; similarly, we denote by $\left(\Sigma_{0}\right)^{\alpha}$ the $\alpha$-upstream component defined as the set of games reached by an $\alpha$-upstream path starting from a game in $\Sigma_{0}$.
@@
For a ‘sophisticated enough model’, it is reasonable (by the construction of Theorem 25 above) to postulate that all the (playable) Xent Games are related in the following sense: one can always modify a playable game into another one by a chain that is $\alpha$-downstream and by a chain that is $\alpha$-upstream, for some small enough $\alpha$. This is the content of the following:
@@
This may suggest to take $\alpha$ small enough to cover the whole space of games (note that if $\alpha$ is too small, games will only be marginally related); or we could also instead decide to focus on an $\alpha$-connected component for some a priori fixed $\Sigma_{0}$. Generally speaking, the choice of $\alpha$ seems to be a subtle question which we plan on investigating in further works.